# Contact graphs of boxes with unidirectional contactsthanks: This research is partially supported by the ANR GATO, under contract ANR-16-CE40-0009 and ANR project GRALMECO (ANR-21-CE48-0004-01)

### v3

@@ 3 Relation with Cover Graphs
The class of CBU graphs is strictly contained in the class of cover graphs. In Lemma 16 we will exhibit a graph that is not CBU but is a cover graph.
@@ 4 Boxicity
Proof. Consider a bipartite graph $G$ with vertex sets $A$ and $B$. Consider a boxicity $b$ representation of $G$ and slightly expand each box in such a way that the intersection graph remains unchanged (we stop the expansion of the boxes before creating new intersections). Now, any two intersecting boxes intersect on a $b$-dimensional box. Assume that this representation is drawn in the space spanned by $e_{2},\ldots,e_{b+1}$, and let us set for the first dimension (spanned by $e_{1}$) that the vertices of $A$ and $B$, correspond to the intervals $[0,1]$ and $[1,2]$, respectively. As $A$ and $B$ are independent sets, it is clear that the boxes in the representation are interior disjoint and that any two intersecting boxes intersect on a $b$-dimensional box orthogonal to $e_{1}$. The obtained representation is thus a $(b+1)$-CBU representation of $G$. $\Box$
Theorem 6 does not extend to triangle-free graphs. We will see in the following section that there exists triangle-free graphs with bounded boxicity that are not $d$-CBU, for any value $d$. Actually, Lemma 16 tells that there exists such graphs with girth 5. In other words, for a $3\leq g\leq 5$, there is no function $f_{g}$ such that every graph $G$, of girth at least $g$ and of boxicity $b$ belongs to $(f_{g}(b))$-CBU.
Problem 7
@@
By Theorem 12, we know that if $f_{6}$ exists, then $f_{6}(2)$ is at least 3. Nevertheless, the following theorem shows that subdividing the edges enables to consider every triangle-free graph. An intersection representation is said proper if two objects intersect if and only if some point of the representation belongs to these 2 objects, only.
Theorem 8
@@
Proof. Consider such a representation of $G$ and slightly expand each box in such a way that the intersection graph remains unchanged, and any two intersecting boxes intersect on a $b$-dimensional box. Assume that this representation is drawn in the space spanned by $e_{2},\ldots,e_{b+1}$, and for the first dimension (spanned by $e_{1}$) let us consider any vertex ordering, $v_{1},\ldots,v_{n}$. For the first dimension, a vertex $v_{i}$, corresponds to the interval $[2i,2i+1]$. Clearly, none of these boxes intersect. Let us now add the boxes for the vertices added by subdividing the edges of $G$. For any edge $v_{i}v_{j}$, in the space spanned by $e_{2},\ldots,e_{b+1}$, the expansion ensured that the intersection of $v_{i}$ and $v_{j}$ contains a box $B_{ij}$, that does not intersect any other box of the representation. If $i<j$, the subdivision vertex of $v_{i}v_{j}$, is represented by $[2i+1,2j]\times B_{i,j}$. The obtained representation is clearly a $(b+1)$-CBU representation of the subdivision of $G$. $\Box$
Corollary 9
@@ 5 Planar graphs
While planar graphs have boxicity at most 3 [35, 18, 4], many subclasses of planar graphs are known to have boxicity at most 2. This is the case for 4-connected planar graphs [34], and their subgraphs. The subgraphs of 4-connected graphs include every triangle-free planar graph (see Lemma 4.1 in [20]). As observed earlier, for those the representation is necessarily proper. For general planar graphs, the representation in $\operatorname{\mathbb{R}}^{3}$ provided in [18] is clearly proper. So Theorem 8 implies the following.
Corollary 10
@@
Proof. Let us prove that for any connected outerplanar graph $G$, and any facial walk $v_{1},v_{2},\ldots,v_{k},v_{k+1}=v_{1}$ on the outerboundary of $G$ (with separating vertices appearing several times in this walk), there exists a 2-CBU representation of $G$ with top sequence $v_{1},v_{2},\ldots,v_{k},v_{k+1}$.
@@
In the first case we can add the box of $v_{i}$ in the representation of $G\setminus v_{i}$ obtained by induction (see Figure 3, left). In the second case we consider the representation of $G\setminus\{v_{i+1},\ldots,v_{j-1}\}$ obtained by induction. Note that $v_{i}$ and $v_{j}$ now appear consecutively in its outerboundary, such as in the top sequence. It is thus easy to add the boxes of $v_{i+1},\ldots,v_{j-1}$ in the representation (see Figure 3, right). $\Box$
@@
Figure 5: The series-parallel graph $G$ of Theorem 12.
Theorem 12
@@
Proof. Consider a graph $G$ with two vertices, $a$ and $b$, linked by 9 disjoint $ab$-paths of length three $ax_{i}y_{i}b$, for $i\in\{1,\ldots,9\}$. Then for each edge $x_{i}y_{i}$ add a length 5 path from $x_{i}$ to $y_{i}$. The obtained graph has girth 6 and is series-parallel (see Figure 5).
Note that in a 2-CBU representation of $G$, the box of $a$ (resp. $b$) has at most four neighbors such that their intersection contains a corner of $a$ (resp. $b$). Thus, there exists an $i\in\{1,\ldots,9\}$ such that one side of $x_{i}$ is contained in one side of $a$, and one side of $y_{i}$ is contained in one side of $b$. Now, whatever the way $x_{i}$ and $y_{i}$ intersect (a side of $x_{i}$ may be contained in a side of $y_{i}$, or the other way around, or also their intersection may contain a corner of each box), it is not possible to have the length 5 $x_{i}y_{i}$-path (see Figure 5, right). If a side of $x_{i}$ is contained in a side of $y_{i}$ there is no place left around $x_{i}$ to draw a third neighbor. If the intersection of $x_{i}$ and $y_{i}$ contains a corner of $x_{i}$ and a corner of $y_{i}$, there is space to draw a third neighbor for these vertices, say $u$ and $v$ respectively, but in that case the $uv$-path should go around $a$ or $b$, but it would intersect the paths $ax_{j}y_{j}b$ with $j\neq i$. Thus $G$ does not admit a 2-CBU representation. $\Box$
Problem 13
@@
Proof. For any 2-CBU representation of the cycle $C$ of length $g$ there is a rectangle $R$, for example the one with the leftmost right side, such that none of the top or bottom side of $R$ is incident to the inner region. It is thus impossible to connect $R$ with a ray in the inner region. On the other hand, there is no planar embedding of $W^{\prime}_{g}$ where $C$ bounds an inner face. $\Box$
@@
Proof. Let us consider the graph $G_{1}$ represented in Figure 7. One can show that this graph does not admit any valid CBU orientation where in the $C_{4}$ induced by $a,b,c$ and $d$, $a$ is a source and $d$ is a sink (nor the converse). Let us assume that there exists a CBU orientation such that $a$ is a source and $d$ is a sink. By fixing the orientation of edges $ab$ and $ac$ from $a$ to $b$ and from $a$ to $c$ respectively, and by Claim 2, the edges $be$ and $cf$ have to be oriented from $b$ to $e$ and from $c$ to $f$. The edge $ef$ is oriented in any direction, w.l.o.g. let us say from $e$ to $f$. But in that case, in the $C_{5}$ induced by $b,e,f,c$ and $d$ contains two sources and two sinks, which is not a valid CBU orientation (By Claim 2).
@@
Then in the partial orientation obtained, we can conclude that in the $C_{4}$ induced by $a,c,i,$ and $j$, the vertex $c$ will be a sink and vertex $j$ will be a source. However, as mentioned in the beginning of this proof, this orientation will not lead to valid orientation, since $j$ and $c$ play the same role as $a$ and $d$ in $G_{1}$. Hence, $G_{3}$ does not admit a valid CBU orientation. $\Box$
Remark 17
The graph $G_{3}$ used in the proof of Lemma 16 is actually a cover graph. In Figure 7, the bottom picture depicts its Hasse diagram.
Since every planar graph with girth at least 10 has circular chromatic number at most $5/2$ [14], the forthcoming Theorem 35 implies that such a graph necessarily belongs to CBU.
Problem 18
@@ 6 Structural properties of $d$-CBU and CBU
Proof. Let us begin with the simpler "only if" part. We proceed as in the proof of Theorem 6 in order to obtain a $(d+1)$-CBU representation of $B$ such that every vertex of $X$ (resp. $Y$) corresponds to $[0,1]$ (resp. $[1,2]$) in the space spanned by $e_{1}$. Then it suffices to add the boxes for $x,y$ and $z$. For a sufficiently large $\Omega$, $x$ is represented by $[-1,0]\times[-\Omega,+\Omega]\times\ldots\times[-\Omega,+\Omega]$, $y$ is represented by $[2,3]\times[-\Omega,+\Omega]\times\ldots\times[-\Omega,+\Omega]$, and $z$ is represented by $[0,2]\times[\Omega-1,\Omega]\times\ldots\times[\Omega-1,\Omega]$.
@@
This orientation ensures that the labels of all the arcs is the same. This implies that there is an hyperplane $\operatorname{\mathcal{H}}$ orthogonal to $e_{1}$ such that for any pair of intersecting boxes $x^{\prime}\in X$ and $y^{\prime}\in Y$, their intersection belongs to $\operatorname{\mathcal{H}}$. This implies that projecting the $(d+1)$-CBU representation (restricted to $B$) along $e_{1}$ leads to a boxicity $d$ representation of $B$. $\Box$
@@
Proof. Let $A,B$ be the parts of $K_{a,b}$. Given a CBU representation of $G$ in $\operatorname{\mathbb{R}}^{d}$ we are going to build a CBU representation of $H$ in $\operatorname{\mathbb{R}}^{d+1}$. For this, the first $d$ intervals defining each $d$-box remain unchanged while the last interval is $[0,1]$ for the vertices in $A$, $[2,3]$ for the vertices in $B$, and $[0,3]$ for the remaining vertices. It is now easy to check that two boxes intersect if and only if they intersect in $G$ and if they are not adjacent in $K_{a,b}$. It is also clear that the intersections occur on planes orthogonal to $e_{1}$. $\Box$
@@
Proof. The "if" part is obvious as $G$ is an induced subgraph of $G^{v}$. For the "only if" part, given a CBU representation of $G$ in $\operatorname{\mathbb{R}}^{d}$ we are going to build a CBU representation of $G^{v}$ in $\operatorname{\mathbb{R}}^{d+1}$. For this, the first $d$ intervals defining each $d$-box remain unchanged, and those of $v^{\prime}$ are the same as those of $v$. The last interval is $[0,1]$ for $v$, $[2,3]$ for $v^{\prime}$, and $[0,3]$ for all the remaining vertices. It is now easy to check that two boxes intersect if and only if they intersected and if one of them is distinct from $v$ or $v^{\prime}$. It is also clear that the intersections occur on planes orthogonal to $e_{1}$. $\Box$
@@
Proof. This clearly holds for the one vertex graph $H_{2}$. By induction on $m$ consider a representation of $H_{m-1}$, add a false twin for every vertex $(i,m-1)$ and modify the first interval of these new twins, so that the interval $[i,m-1]$ becomes $[i,m]$. These boxes correspond to the vertices $(i,m)$ with $i<m-1$. For the vertex $(m-1,m)$, one should add a box $[m-1,m]\times[-\Omega,+\Omega]\times\ldots\times[-\Omega,+\Omega]$, for a sufficiently large $\Omega$. To deal with the intersections between this box and the boxes of the other vertices $(i,m)$, we add a new dimension such that vertex $(m-1,m)$ has interval $[1,2]$, the vertices $(i,m-1)$ have interval $[1,2]$, the vertices $(i,m)$ with $i<m-1$ have interval $[3,4]$, and all the other vertices have interval $[1,4]$. $\Box$
Theorem 24
@@
Proof. We have already seen that $a)\Rightarrow b)$. Let us show $b)\Rightarrow c)$. Consider a homogeneous arc labeling of $G$, with labels in $[2,m-1]$, for the minimum $m$. By minimality of $m$, note that all the labels are used, and thus $m-2\leq n-1$. Let $H_{m}^{t}$ be the graph obtained from the shift graph $H_{m}$ by adding $t_{i,j}$ false twins of vertex $(i,j)$ if there are $t_{i,j}+1$ vertices of $G$ whose incoming arcs are labeled $i$, and whose outgoing arcs are labeled $j$. For the vertices without incoming (resp. outgoing) arcs assume that those are labeled 1 (resp. $m$). Consider now an injective mapping $\gamma\ :\ V(G)\longrightarrow V(H_{m}^{t})$, such that any vertex with incoming and outgoing arcs labeled $i,j$ is mapped to $(i,j)$ or one of its twins. This mapping ensures us that $G$ is a subgraph of $H_{m}^{t}$. Indeed, for any two adjacent vertices $u,v$ of $G$ linked by an edge labelled $j$ oriented from $u$ to $v$, their incoming and outgoing arcs are labeled $i,j$ and $j,k$ respectively, for some $i<j<k$, and thus the vertices $\gamma(u)$ and $\gamma(v)$ of $H_{m}^{t}$ are adjacent, as they correspond to or are twins of $(i,j)$ and $(j,k)$.
We now show $c)\Rightarrow d)$. Consider a graph $H_{m}^{t}$ containing $G$ as a subgraph, for some $m,t$ such that $m+t\leq n+1$. By Theorem 23 and Theorem 22 we have that $H_{m}^{t}$ belongs to $(m-1+t)$-CBU, and so to $n$-CBU. Starting from $H_{m}^{t}$ one can obtain $G$ by successively deleting $n-1$ stars $K_{1,b}$, so by Theorem 21, we have that $G$ belongs to $(2n-1)$-CBU. Finally, $d)\Rightarrow a)$ is obvious. $\Box$
It is easy to see that every complete bipartite graph belongs to 3-CBU. By Theorem 21, removing stars $K_{1,b}$ centered on the smallest part, one obtains that every $n$-vertex bipartite graph belongs to $(\lfloor n/2\rfloor+3)$-CBU. One can reach a slightly better bound from Theorem 6, and the fact that for every graph $G$, $box(G)\leq\lfloor n/2\rfloor$ [31].
Corollary 25
@@
Proof. Indeed, for any $g\geq 3$ there exist graphs of girth $g$ with fractional chromatic number at least 4 [16]. (Actually, their fractional chromatic number is arbitrarily large). By Theorem 39, such graphs cannot belong to CBU. $\Box$ Nevertheless, the following remains open.
Problem 27

### v4

@@ 3 Relation with Cover Graphs
The class of CBU graphs is strictly contained in the class of cover graphs. In Lemma 5.21 we will exhibit a graph that is not CBU but is a cover graph.
@@ 4 Boxicity
Proof 4.7.
Consider a bipartite graph $G$ with vertex sets $A$ and $B$. Consider a boxicity $b$ representation of $G$ and slightly expand each box in such a way that the intersection graph remains unchanged (we stop the expansion of the boxes before creating new intersections). Now, any two intersecting boxes intersect on a $b$-dimensional box. Assume that this representation is drawn in the space spanned by $e_{2},\ldots,e_{b+1}$, and let us set for the first dimension (spanned by $e_{1}$) that the vertices of $A$ and $B$, correspond to the intervals $[0,1]$ and $[1,2]$, respectively. As $A$ and $B$ are independent sets, it is clear that the boxes in the representation are interior disjoint and that any two intersecting boxes intersect on a $b$-dimensional box orthogonal to $e_{1}$. The obtained representation is thus a $(b+1)$-CBU representation of $G$.
Theorem 6 does not extend to triangle-free graphs. We will see in the following section that there exists triangle-free graphs with bounded boxicity that are not $d$-CBU, for any value $d$. Actually, Lemma 5.21 tells that there exists such graphs with girth 5. In other words, for a $3\leq g\leq 5$, there is no function $f_{g}$ such that every graph $G$, of girth at least $g$ and of boxicity $b$ belongs to $(f_{g}(b))$-CBU.
Problem 4.8.
@@
By Theorem 5.15, we know that if $f_{6}$ exists, then $f_{6}(2)$ is at least 3. Nevertheless, the following theorem shows that subdividing the edges enables to consider every triangle-free graph. An intersection representation is said proper if two objects intersect if and only if some point of the representation belongs to these 2 objects, only.
Theorem 4.9.
@@
Proof 4.10.
Consider such a representation of $G$ and slightly expand each box in such a way that the intersection graph remains unchanged, and any two intersecting boxes intersect on a $b$-dimensional box. Assume that this representation is drawn in the space spanned by $e_{2},\ldots,e_{b+1}$, and for the first dimension (spanned by $e_{1}$) let us consider any vertex ordering, $v_{1},\ldots,v_{n}$. For the first dimension, a vertex $v_{i}$, corresponds to the interval $[2i,2i+1]$. Clearly, none of these boxes intersect. Let us now add the boxes for the vertices added by subdividing the edges of $G$. For any edge $v_{i}v_{j}$, in the space spanned by $e_{2},\ldots,e_{b+1}$, the expansion ensured that the intersection of $v_{i}$ and $v_{j}$ contains a box $B_{ij}$, that does not intersect any other box of the representation. If $i<j$, the subdivision vertex of $v_{i}v_{j}$, is represented by $[2i+1,2j]\times B_{i,j}$. The obtained representation is clearly a $(b+1)$-CBU representation of the subdivision of $G$.
Corollary 4.11.
@@ 5 Planar graphs
While planar graphs have boxicity at most 3 [35, 18, 4], many subclasses of planar graphs are known to have boxicity at most 2. This is the case for 4-connected planar graphs [34], and their subgraphs. The subgraphs of 4-connected graphs include every triangle-free planar graph (see Lemma 4.1 in [20]). As observed earlier, for those the representation is necessarily proper. For general planar graphs, the representation in $\operatorname{\mathbb{R}}^{3}$ provided in [18] is clearly proper. So Theorem 4.9 implies the following.
Corollary 5.12.
@@
Proof 5.14.
Let us prove that for any connected outerplanar graph $G$, and any facial walk $v_{1},v_{2},\ldots,v_{k},v_{k+1}=v_{1}$ on the outerboundary of $G$ (with separating vertices appearing several times in this walk), there exists a 2-CBU representation of $G$ with top sequence $v_{1},v_{2},\ldots,v_{k},v_{k+1}$.
@@
In the first case we can add the box of $v_{i}$ in the representation of $G\setminus v_{i}$ obtained by induction (see Figure 3, left). In the second case we consider the representation of $G\setminus\{v_{i+1},\ldots,v_{j-1}\}$ obtained by induction. Note that $v_{i}$ and $v_{j}$ now appear consecutively in its outerboundary, such as in the top sequence. It is thus easy to add the boxes of $v_{i+1},\ldots,v_{j-1}$ in the representation (see Figure 3, right).
@@
Figure 5: The series-parallel graph $G$ of Theorem 5.15.
Theorem 5.15.
@@
Proof 5.16.
Consider a graph $G$ with two vertices, $a$ and $b$, linked by 9 disjoint $ab$-paths of length three $ax_{i}y_{i}b$, for $i\in\{1,\ldots,9\}$. Then for each edge $x_{i}y_{i}$ add a length 5 path from $x_{i}$ to $y_{i}$. The obtained graph has girth 6 and is series-parallel (see Figure 5).
Note that in a 2-CBU representation of $G$, the box of $a$ (resp. $b$) has at most four neighbors such that their intersection contains a corner of $a$ (resp. $b$). Thus, there exists an $i\in\{1,\ldots,9\}$ such that one side of $x_{i}$ is contained in one side of $a$, and one side of $y_{i}$ is contained in one side of $b$. Now, whatever the way $x_{i}$ and $y_{i}$ intersect (a side of $x_{i}$ may be contained in a side of $y_{i}$, or the other way around, or also their intersection may contain a corner of each box), it is not possible to have the length 5 $x_{i}y_{i}$-path (see Figure 5, right). If a side of $x_{i}$ is contained in a side of $y_{i}$ there is no place left around $x_{i}$ to draw a third neighbor. If the intersection of $x_{i}$ and $y_{i}$ contains a corner of $x_{i}$ and a corner of $y_{i}$, there is space to draw a third neighbor for these vertices, say $u$ and $v$ respectively, but in that case the $uv$-path should go around $a$ or $b$, but it would intersect the paths $ax_{j}y_{j}b$ with $j\neq i$. Thus $G$ does not admit a 2-CBU representation.
Problem 5.17.
@@
Proof 5.19.
For any 2-CBU representation of the cycle $C$ of length $g$ there is a rectangle $R$, for example the one with the leftmost right side, such that none of the top or bottom side of $R$ is incident to the inner region. It is thus impossible to connect $R$ with a ray in the inner region. On the other hand, there is no planar embedding of $W^{\prime}_{g}$ where $C$ bounds an inner face.
@@
Proof 5.22.
Let us consider the graph $G_{1}$ represented in Figure 7. One can show that this graph does not admit any valid CBU orientation where in the $C_{4}$ induced by $a,b,c$ and $d$, $a$ is a source and $d$ is a sink (nor the converse). Let us assume that there exists a CBU orientation such that $a$ is a source and $d$ is a sink. By fixing the orientation of edges $ab$ and $ac$ from $a$ to $b$ and from $a$ to $c$ respectively, and by Claim 2, the edges $be$ and $cf$ have to be oriented from $b$ to $e$ and from $c$ to $f$. The edge $ef$ is oriented in any direction, w.l.o.g. let us say from $e$ to $f$. But in that case, in the $C_{5}$ induced by $b,e,f,c$ and $d$ contains two sources and two sinks, which is not a valid CBU orientation (By Claim 2).
@@
Then in the partial orientation obtained, we can conclude that in the $C_{4}$ induced by $a,c,i,$ and $j$, the vertex $c$ will be a sink and vertex $j$ will be a source. However, as mentioned in the beginning of this proof, this orientation will not lead to valid orientation, since $j$ and $c$ play the same role as $a$ and $d$ in $G_{1}$. Hence, $G_{3}$ does not admit a valid CBU orientation.
Remark 5.23.
The graph $G_{3}$ used in the proof of Lemma 5.21 is actually a cover graph. In Figure 7, the bottom picture depicts its Hasse diagram.
Since every planar graph with girth at least 10 has circular chromatic number at most $5/2$ [14], the forthcoming Theorem 7.48 implies that such a graph necessarily belongs to CBU.
Problem 5.24.
@@ 6 Structural properties of $d$-CBU and CBU
Proof 6.27.
Let us begin with the simpler "only if" part. We proceed as in the proof of Theorem 6 in order to obtain a $(d+1)$-CBU representation of $B$ such that every vertex of $X$ (resp. $Y$) corresponds to $[0,1]$ (resp. $[1,2]$) in the space spanned by $e_{1}$. Then it suffices to add the boxes for $x,y$ and $z$. For a sufficiently large $\Omega$, $x$ is represented by $[-1,0]\times[-\Omega,+\Omega]\times\ldots\times[-\Omega,+\Omega]$, $y$ is represented by $[2,3]\times[-\Omega,+\Omega]\times\ldots\times[-\Omega,+\Omega]$, and $z$ is represented by $[0,2]\times[\Omega-1,\Omega]\times\ldots\times[\Omega-1,\Omega]$.
@@
This orientation ensures that the labels of all the arcs is the same. This implies that there is an hyperplane $\operatorname{\mathcal{H}}$ orthogonal to $e_{1}$ such that for any pair of intersecting boxes $x^{\prime}\in X$ and $y^{\prime}\in Y$, their intersection belongs to $\operatorname{\mathcal{H}}$. This implies that projecting the $(d+1)$-CBU representation (restricted to $B$) along $e_{1}$ leads to a boxicity $d$ representation of $B$.
@@
Proof 6.29.
Let $A,B$ be the parts of $K_{a,b}$. Given a CBU representation of $G$ in $\operatorname{\mathbb{R}}^{d}$ we are going to build a CBU representation of $H$ in $\operatorname{\mathbb{R}}^{d+1}$. For this, the first $d$ intervals defining each $d$-box remain unchanged while the last interval is $[0,1]$ for the vertices in $A$, $[2,3]$ for the vertices in $B$, and $[0,3]$ for the remaining vertices. It is now easy to check that two boxes intersect if and only if they intersect in $G$ and if they are not adjacent in $K_{a,b}$. It is also clear that the intersections occur on planes orthogonal to $e_{1}$.
@@
Proof 6.31.
The "if" part is obvious as $G$ is an induced subgraph of $G^{v}$. For the "only if" part, given a CBU representation of $G$ in $\operatorname{\mathbb{R}}^{d}$ we are going to build a CBU representation of $G^{v}$ in $\operatorname{\mathbb{R}}^{d+1}$. For this, the first $d$ intervals defining each $d$-box remain unchanged, and those of $v^{\prime}$ are the same as those of $v$. The last interval is $[0,1]$ for $v$, $[2,3]$ for $v^{\prime}$, and $[0,3]$ for all the remaining vertices. It is now easy to check that two boxes intersect if and only if they intersected and if one of them is distinct from $v$ or $v^{\prime}$. It is also clear that the intersections occur on planes orthogonal to $e_{1}$.
@@
Proof 6.33.
This clearly holds for the one vertex graph $H_{2}$. By induction on $m$ consider a representation of $H_{m-1}$, add a false twin for every vertex $(i,m-1)$ and modify the first interval of these new twins, so that the interval $[i,m-1]$ becomes $[i,m]$. These boxes correspond to the vertices $(i,m)$ with $i<m-1$. For the vertex $(m-1,m)$, one should add a box $[m-1,m]\times[-\Omega,+\Omega]\times\ldots\times[-\Omega,+\Omega]$, for a sufficiently large $\Omega$. To deal with the intersections between this box and the boxes of the other vertices $(i,m)$, we add a new dimension such that vertex $(m-1,m)$ has interval $[1,2]$, the vertices $(i,m-1)$ have interval $[1,2]$, the vertices $(i,m)$ with $i<m-1$ have interval $[3,4]$, and all the other vertices have interval $[1,4]$.
Theorem 6.34.
@@
Proof 6.35.
We have already seen that $a)\Rightarrow b)$. Let us show $b)\Rightarrow c)$. Consider a homogeneous arc labeling of $G$, with labels in $[2,m-1]$, for the minimum $m$. By minimality of $m$, note that all the labels are used, and thus $m-2\leq n-1$. Let $H_{m}^{t}$ be the graph obtained from the shift graph $H_{m}$ by adding $t_{i,j}$ false twins of vertex $(i,j)$ if there are $t_{i,j}+1$ vertices of $G$ whose incoming arcs are labeled $i$, and whose outgoing arcs are labeled $j$. For the vertices without incoming (resp. outgoing) arcs assume that those are labeled 1 (resp. $m$). Consider now an injective mapping $\gamma\ :\ V(G)\longrightarrow V(H_{m}^{t})$, such that any vertex with incoming and outgoing arcs labeled $i,j$ is mapped to $(i,j)$ or one of its twins. This mapping ensures us that $G$ is a subgraph of $H_{m}^{t}$. Indeed, for any two adjacent vertices $u,v$ of $G$ linked by an edge labelled $j$ oriented from $u$ to $v$, their incoming and outgoing arcs are labeled $i,j$ and $j,k$ respectively, for some $i<j<k$, and thus the vertices $\gamma(u)$ and $\gamma(v)$ of $H_{m}^{t}$ are adjacent, as they correspond to or are twins of $(i,j)$ and $(j,k)$.
We now show $c)\Rightarrow d)$. Consider a graph $H_{m}^{t}$ containing $G$ as a subgraph, for some $m,t$ such that $m+t\leq n+1$. By Theorem 6.32 and Theorem 6.30 we have that $H_{m}^{t}$ belongs to $(m-1+t)$-CBU, and so to $n$-CBU. Starting from $H_{m}^{t}$ one can obtain $G$ by successively deleting $n-1$ stars $K_{1,b}$, so by Theorem 6.28, we have that $G$ belongs to $(2n-1)$-CBU. Finally, $d)\Rightarrow a)$ is obvious.
It is easy to see that every complete bipartite graph belongs to 3-CBU. By Theorem 6.28, removing stars $K_{1,b}$ centered on the smallest part, one obtains that every $n$-vertex bipartite graph belongs to $(\lfloor n/2\rfloor+3)$-CBU. One can reach a slightly better bound from Theorem 6, and the fact that for every graph $G$, $box(G)\leq\lfloor n/2\rfloor$ [31].
Corollary 6.36.
@@
Proof 6.38.
Indeed, for any $g\geq 3$ there exist graphs of girth $g$ with fractional chromatic number at least 4 [16]. (Actually, their fractional chromatic number is arbitrarily large). By Theorem 8.54, such graphs cannot belong to CBU.
Nevertheless, the following remains open.
Problem 6.39.